OpenStudy (anonymous):

f(x-1) where f(x) = (x+1)^5 -3x^2

OpenStudy (mertsj):

\[f(x-1)=(x-1+1)^5-3(x-1)^2=x^5-3(x^2-2x+1)=\]

OpenStudy (mertsj):

\[x^5-3x^2+6x-3\]

OpenStudy (anonymous):

Thank you